### v1

# CRAFT-LoRA: Content-Style Personalization via Rank-Constrained Adaptation and Training-Free Fusion

Abstract
Personalized image generation requires effectively balancing content fidelity with stylistic consistency when synthesizing images based on text and reference examples. Low-Rank Adaptation (LoRA) offers an efficient personalization approach, with potential for precise control through combining LoRA weights on different concepts. However, existing combination techniques face persistent challenges: entanglement between content and style representations, insufficient guidance for controlling elements’ influence, and unstable weight fusion that often require additional training. We address these limitations through CRAFT-LoRA, with complementary components: (1) rank-constrained backbone fine-tuning that injects low-rank projection residuals to encourage learning decoupled content and style subspaces; (2) a prompt-guided approach featuring an expert encoder with specialized branches that enables semantic extension and precise control through selective adapter aggregation; and (3) a training-free, timestep-dependent classifier-free guidance scheme that enhances generation stability by strategically adjusting noise predictions across diffusion steps. Our method significantly improves content-style disentanglement, enables flexible semantic control over LoRA module combinations, and achieves high-fidelity generation without additional retraining overhead.
Figure 1: Samples generated by our proposed framework. Our method achieves more effective decoupling and fusion of content and style, enabling finer control over both aspects during generation.
11footnotetext: This work was done as a visiting student at AGI Lab.22footnotetext: Corresponding author.
1 Introduction
Recent advancements in deep learning have fueled remarkable progress in generative models, particularly within the domain of personalized image generation [11, 2, 30]. Unlike generic image synthesis, personalized generation focuses on customizing outputs based on user-provided references, such as a specific subject’s appearance or stylistic attributes. In more sophisticated scenarios, these systems aim to simultaneously preserve subject identity while emulating a target style, enabling applications in creative design, entertainment, digital avatars, personalized marketing, and beyond.
Among the various techniques developed for personalized image generation, Low-Rank Adaptation (LoRA) has emerged as a particularly prominent and efficient approach [15]. LoRA facilitates adaptation by introducing small, trainable low-rank matrices into specific layers of large pre-trained text-to-image diffusion models, thereby allowing for efficient fine-tuning on new concepts using only a small number of references. When integrated with personalization frameworks like DreamBooth, LoRA-based methods enable the customization of generated images to specific content, e.g. preserving the identity of a person, or style, e.g. emulating a particular artistic rendering or texture, making them widely adopted [29]. Moreover, LoRA exhibits a unique compositional property: multiple independently trained LoRA modules, each corresponding to different visual concepts, can be simultaneously applied during inference. This enables the model to generate images that blend diverse attributes, such as combining a specific subject with a distinct artistic style. However, the combinatorial use of LoRA weights introduces new challenges. Naively merging multiple LoRA modules often leads to degraded visual quality or semantic inconsistency, due to entangled or conflicting latent representations. Consequently, recent works have explored various strategies for more effective LoRA fusion, including approaches that re-optimize weights for specific combinations [31], adjust internal attention mechanisms [9], or dynamically modulate different LoRAs’ influence during generation [22].
Despite recent progress, existing LoRA combination strategies still face several fundamental challenges, which motivate the improvements proposed in this paper. First, although pre-trained diffusion models empirically exhibit some degree of composability with respect to LoRA modules, these models are not explicitly trained to support such combinations. As a result, directly merging LoRA weights often fails to achieve a clean disentanglement between content and style factors present in the reference images [19, 18, 1]. Inspired by the meta-learning paradigm, particularly Model-Agnostic Meta-Learning (MAML) [8, 6], which seeks optimal initializations for fast and generalizable task adaptation, we propose a novel network re-initialization strategy for pre-trained diffusion models that enhances their compatibility with compositional LoRA usage. Specifically, we introduce a low-rank constrained adaptation phase over the parameters of the pre-trained diffusion model inspired by PaRa [5]. The additional training stage narrows the model’s generative space, shifting it from a broadly generalized prior to a more compact and disentangled representation space that better supports the fusion of multiple LoRA modules.
Another critical limitation lies in the inherent granularity ambiguity associated with the definition of a “subject” in personalized image generation. In practical scenarios, the notion of a subject often encompasses a hierarchy of visual attributes—for instance, a portrait image may include multiple levels of subject-specific features, e.g. facial identity, hairstyle, clothing style, and accessory details. However, current methods typically collapse this rich semantic structure into a single coarse-grained token representation within the input prompt [24, 20]. This oversimplification neglects the internal structure of subject attributes and lacks the mechanisms to differentiate or control the preservation and variation of fine-grained elements.
To address this, we propose a novel expert encoder system that enables fine-grained control over subject features at different levels. Our system implements specially designed content and style adapters operating in disjoint network layers. The expert encoder processes identity labels, content descriptions, and style specifications through separate branches, facilitating selective weights aggregation based on semantic relevance The designed expert system enables precise delineation of feature preservation hierarchies through prompts, allowing for discriminative activation of specific subject attributes at variable granularity levels according to specified semantic constraints.
Finally, existing LoRA fusion strategies, such as ZipLoRA [31], typically rely on additional optimization procedures to reconcile differences between LoRA modules [37, 12]. While effective to some extent, directly modifying LoRA parameters during fusion can inadvertently alter or suppress critical elements encoded in the original weights, leading to loss of identity or stylistic fidelity. Moreover, these methods incur non-trivial computational overhead. To overcome these limitations, we propose a novel tuning-free LoRA combination method that circumvents the need for parameter-level fusion. At the core of our approach is our developed asymmetric Classifier-Free Guidance (CFG) mechanism applied during the diffusion process. Rather than adjusting LoRA weights directly, we selectively apply time-dependent modifications only to the conditional path where combined LoRA weights are applied, while keeping the unconditional path that uses only the base model without LoRA adapters fixed to the base model weights. Our asymmetric design yields a dynamic guidance signal that effectively steers the generation process toward a desirable balance between content and style, without introducing additional training cost or risking the degradation of essential features, thereby offering a lightweight and robust alternative to existing LoRA fusion techniques.
As illustrated in Figure 2, our proposed unified framework integrates rank-constrained fine-tuning to learn decoupled content and style subspaces, leverages prompt-guided mechanisms for semantic control and extension via selective adapter aggregation guided by an expert encoder, and employs a time-step-aware guidance correction scheme for stable, high-fidelity generation without extra training.
To summarize, we make the following key contributions:
  • We propose a novel framework that enhances content-style disentanglement during LoRA training through rank-constrained fine-tuning and low-rank projection residuals, encouraging the learning of decoupled subspaces.
  • We introduce a prompt-guided approach with an expert encoder and selective adapter aggregation, enabling precise semantic control over content and style influences and extending the applicability of LoRA modules.
  • We develop a training-free time-step-aware and classifier-free guidance correction scheme that improves the stability and fidelity of diffusion-based generation by strategically adjusting noise predictions during critical steps.
Figure 2: Overview of CRAFT-LoRA, a unified pipeline for personalized image synthesis. In the Training Stage, content ($\Delta W_{c}$) and style ($\Delta W_{s}$) LoRA adapters are decoupled from reference images using a rank-restricted initialization offset. The Prompt Text Guidance module employs an expert system with specialized branches to produce distinct content and style tokens for semantic control. In the Inference Stage, a timestep-aware asymmetric CFG scheme selectively integrates LoRA updates, ensuring stable and high-fidelity image generation.
2 Related Work
2.1 Personalized image generation
The research trajectory of image stylization has evolved from early Generative Adversarial Network-based domain translation methods—often relying on paired or unpaired datasets and extensive, domain-specific training—to diffusion-based frameworks [4, 26, 3]. More recently, the application of large pre-trained vision–language models has further advanced zero-shot editing and stylization. For instance, Prompt-to-Prompt [14] guides the generation process by selectively modulating cross-attention maps, while plug-and-play methods introduce lightweight adapters into self-attention layers to adjust spatial features dynamically [21, 13, 38]. Despite significant progress, reconciling high content fidelity with flexible style injection remains an open challenge [7, 16]. Practice shows that naive parameter fusion strategies, such as weighted averaging, often lead to unintended mixing of content and style features [32]. Consequently, current stylization pipelines often struggle to integrate diverse visual concepts without compromising either semantic coherence or stylistic integrity  [35, 23].
2.2 Stylization through Fine-Tuning Methods
The rapid advancement of text-to-image diffusion models has spurred diverse personalization strategies. Full-model fine-tuning approaches, exemplified by DreamBooth [27], can achieve high-fidelity style transfer but incur substantial computational costs and require lengthy retraining on domain-specific examples. To mitigate these demands, Parameter-Efficient Fine-Tuning techniques have emerged [28]. Among these, LoRA introduces small and trainable low-rank decomposition matrices, injecting new stylistic or semantic concepts while keeping the majority of the pre-trained model weights frozen, thus striking a balance between adaptability and efficiency. Complementing weight-space adaptation methods, token-based optimization techniques, such as Textual Inversion [10], learn new embeddings to represent novel concepts within the model’s token vocabulary. Adapter-based schemes offer an alternative pathway, for example, StyleDrop [33] appends lightweight layers to attention blocks, isolating style information in dedicated modules. Meta-learning, as a cross-task method for domain transfer, is similar in principle to style transfer in personalized image generation, providing new inspirations for fine-tuning methods [39].
2.3 Content-Style LoRAs Merging
Methods for combining multiple LoRA modules, effective for capturing distinct concepts, are explored to achieve simultaneous content and style adaptation. Parameter interpolation is a basic approach, which is prone to suboptimal outcomes, especially when parameters conflict [34, 36]. To address these challenges, several studies focus on refining the composition mechanism or the application process itself. ZipLoRA [31] learns column-wise mixing coefficients to better balance style and content elements, although it introduces the need for additional optimization procedures for each new content-style pairing. B-LoRA [9], in contrast, differentiates the functional roles within attention modules, aiming for a more effective separation of object and style representations. LoRA Composition [40] employs a cyclic update mechanism, allowing LoRA weights to iteratively guide the generation without explicit parameter merging.
3 Methodology
3.1 Preliminaries
LoRA is an efficient fine-tuning method that adapts pretrained diffusion models by adding trainable low-rank matrices $A\in\mathbb{R}^{r\times n}$ and $B\in\mathbb{R}^{m\times r}$ to frozen weights $W_{0}\in\mathbb{R}^{m\times n}$, yielding $W=W_{0}+BA$ with $r\ll\min(m,n)$. In personalized image generation, content LoRAs $\Delta W_{c}$ represent subject-specific features like identity and structure, while style LoRAs $\Delta W_{s}$ capture artistic aspects such as texture and color. Both can be trained efficiently from a single reference image, and during inference they can be composed into the backbone to form personalized generative elements.
3.2 Framework Overview
We propose CRAFT-LoRA, a unified framework for personalized image synthesis that addresses the challenges of content-style disentanglement and fusion in LoRA-based methods. As shown in Fig. 2, the framework includes three main stages: (1) Training Stage, where content and style LoRA weights are learned separately based on a rank-constrained fine-tuned backbone; (2) Prompt Text Guidance, which employs an expert encoder system with specialized branches to provide semantic control over content and style during training and inference; (3) Inference Stage, utilizing a timestep-aware asymmetric classifier-free guidance scheme for stable and flexible fusion of LoRA weights without extra training.
3.3 Rank-Limited Fine-Tuning of Backbone
As mentioned in the introduction, pre-trained diffusion models do not explicitly support clear disentanglement of content and style factors when adapting weights from a reference image, thus separation enhancement requires additional fine-tuning of the diffusion backbone. Motivated by the adaptability of MAML [8, 17] under few-shot constraints and leveraging the rich hierarchical features of SDXL [25], we propose to fine-tune the U-Net backbone under a rank-constrained mechanism. Our approach provides optimized initialization and structurally improves the disentanglement of content and style in the U-Net architecture, providing a more effective foundation for subsequent LoRA separation and personalized synthesis.
3.3.1 Rank Reduction Fine-Tuning.
Inspired by PaRa method [5], we denote the frozen backbone weight matrix of layer $l$ as $W_{l}^{(0)}\in\mathbb{R}^{d_{\mathrm{out}}^{l}\times d_{\mathrm{in}}^{l}}$, with $L$ layers in total. We introduce a learnable basis matrix $B_{l}\in\mathbb{R}^{d_{\mathrm{in}}^{l}\times r_{l}}$ per layer, where $r_{l}$ is the rank constraint satisfying $r_{l}\ll\min(d_{\mathrm{out}}^{l},d_{\mathrm{in}}^{l})$. Computing the QR decomposition $B_{l}=Q_{l}R_{l}$ yields an orthonormal basis $Q_{l}$ with $Q_{l}^{\top}Q_{l}=I$. The rank-limited fine-tuning projects out components of $W_{l}^{(0)}$ along $Q_{l}$, yielding updated weights:
$W_{l}=W_{l}^{(0)}-Q_{l}Q_{l}^{\top}W_{l}^{(0)}.$ (1)
This constrains updates to directions orthogonal to the learned low-rank subspace, introducing an architectural bias that encourages reduced overlap between content and style representations. We show empirically in Appendix. Appendix that it reduces cross-influence.
To further allocate adaptation capacity hierarchically, we schedule:
$r_{l}=r_{\max}-\frac{l-1}{L-1}(r_{\max}-r_{\min}),\quad l=1,\ldots,L,$ (2)
assigning higher ranks to earlier layers, which typically encode low-level structural and textural information where content and style factors are more intertwined, thus requiring more adaptation capacity.
For content-style disentanglement, we train separate basis matrices $B_{l}^{\mathrm{content}}$ and $B_{l}^{\mathrm{style}}$, obtaining corresponding orthonormal subspaces $Q_{l}^{\mathrm{content}}$ and $Q_{l}^{\mathrm{style}}$. Merging these via concatenation and QR decomposition and compute the combined weight update:
$Q_{l}^{\mathrm{merged}}=[Q_{l}^{\mathrm{content}}\mid Q_{l}^{\mathrm{style}}]=Q_{l}^{\mathrm{combined}}R_{l}^{\mathrm{combined}}$ (3)
$W_{l}^{\mathrm{combined}}=W_{l}^{(0)}-Q_{l}^{\mathrm{combined}}(Q_{l}^{\mathrm{combined}})^{\top}W_{l}^{(0)}$ (4)
The fine-tuned backbone weights are retained to provide a targeted initialization, thereby empirically reducing the cross-influence between content and style factors. After rank-limited fine-tuning, we denote the resulting backbone as $W_{\text{init}}=\{W^{\text{init}}_{i}\}_{i=1}^{L}$, which serves as the host weights for all subsequent LoRA training and inference stages.
3.3.2 Contrastive Content-Style Pairs.
To facilitate effective disentanglement of content and style during fine-tuning, we construct a carefully curated dataset consisting of $\mathbf{100}$ contrastive content-style image pairs. These pairs are specifically designed so that either the content or the style remains constant while the other varies, explicitly guiding the model to differentiate and separately capture content and style characteristics.
We operationalize the notion of content and style using frequency-domain decomposition: content corresponds primarily to low-frequency components encoding structural and semantic information such as object layouts, while style corresponds to high-frequency components capturing textures and color palettes. Specifically, we apply Gaussian low-pass filtering with cutoff frequency $\sigma=0.35$ (normalized) to extract content features, and compute the residual as style features. These frequency-separated representations are then used as conditioning signals during text-to-image generation to produce the contrastive pairs.
The content and style samples under this frequency method are shown in Figure 3.
Concretely, we select $\mathbf{10}$ distinct content reference images $\{C_{i}\}_{i=1}^{10}$ and $\mathbf{10}$ distinct style reference images $\{S_{j}\}_{j=1}^{10}$. Exhaustively pairing them yields:
$\{(C_{i},S_{j})\mid i=1,\dots,10;\;j=1,\dots,10\}.$ (5)
Fixing a content image $C_{1}$ and combining it with styles $\{S_{1},\dots,S_{10}\}$ produces 10 pairs that vary only in style, while fixing a style and varying content produces complementary pairs. These contrastive pairs are directly used to drive the rank-limited backbone training described above. For each pair, the two images are fed as separate inputs: one provides supervision for learning the content-specific basis $B_{l}^{\mathrm{content}}$, while the other supervises the style-specific basis $B_{l}^{\mathrm{style}}$. The corresponding orthonormal subspaces $Q_{l}^{\mathrm{content}}$ and $Q_{l}^{\mathrm{style}}$ are then merged, and the backbone is updated according to Eqs.3,4. In this way, the model allocates content and style variations to disjoint low-rank subspaces, ensuring that the backbone initialization $W_{l}^{\mathrm{combined}}$ encodes reduced interference between the two factors.
Empirically, combining rank-limited updates with contrastive pairs leads to a 12% reduction in average CLIP-based content–style similarity, indicating that the learned features exhibit reduced cross-influence. As further demonstrated in Table 6(b), the removal of rank-limited backbone fine-tuning results in clear drops in both content and style fidelity, quantitatively confirming its role in disentanglement. Details on dataset construction and training settings are provided in Appendix Appendix.
Figure 3: Separating content and style components in a single image through frequency domain decomposition
Algorithm 1 Prompt-Guided Decoupling & Aggregation
1:backbone $W_{\text{init}}$; embeddings $e_{\mathrm{sem}},e_{c},e_{s}$; content/style scalars $\gamma_{c},\gamma_{s}$
2:aggregated backbone $W^{\mathrm{agg}}$
3:Initialize updates $\Delta W_{i}^{(c)}\leftarrow 0,\Delta W_{i}^{(s)}\leftarrow 0$ for all $i$
4:for $i=1\to L$ do
5:  if $i\in I_{c}$ then update $\Delta W_{i}^{(c)}\leftarrow B_{i}A_{i}^{(c)}(e_{\mathrm{sem}})$
6:  else if $i\in I_{s}$ then update $\Delta W_{i}^{(s)}\leftarrow B_{i}A_{i}^{(s)}(e_{\mathrm{sem}})$
7:  end if
8:end for
9:Compute $W^{\mathrm{agg}}$ as in Eq. (8)
10:return $W^{\mathrm{agg}}$
Algorithm 2 Asymmetric CFG (ACFG) Sampling
1:backbone $W_{\text{init}}$; LoRA updates $\{\Delta W_{i}^{(c)}\},\{\Delta W_{i}^{(s)}\}$; schedules $T_{c},T_{s}$; guidance $\omega$
2:final sample $\hat{x}_{0}$
3:Initialize $x_{T}\sim\mathcal{N}(0,I)$; for $t=T\to 1$ do
[…]
5:Build $W^{\mathrm{cond}}(t)$ as in Eq. (17); set $W^{\mathrm{uncond}}\leftarrow W_{\text{init}}$
6:Compute $\epsilon_{\mathrm{cond}},\epsilon_{\mathrm{uncond}}$; form $\epsilon_{\mathrm{acfg}}$ as in Eq. (19)
[…]
8:end for
9:return $\hat{x}_{0}$
3.4 Specialized Decoupling and Fusion
Building on the rank-limited backbone initialization, we introduce a prompt-guided framework to further separate content and style at the adapter level. During training, content and style LoRA updates are learned on disjoint layer subsets with explicit prompt markers. At inference, an expert encoder routes prompt information to selectively activate these adapters, enabling flexible and fine-grained fusion.
3.4.1 Prompt-Guided Decoupling and Encoding.
Given a raw prompt $p$ containing explicit markers for content and style, e.g. p=p=”A photo of a person <c>smiling in a watercolor style <s>”, we treat these markers as routing signals to regulate LoRA module activation and to prevent leakage of content semantics into the style branch and vice versa. The prompt with markers removed is encoded by the SDXL text encoder to produce a unified semantic embedding:
$e_{\mathrm{sem}}=\mathrm{SDXL\_Encoder}\bigl(p\setminus\{\texttt{<c>},\texttt{</c>},\texttt{<s>},\texttt{</s>}\}\bigr),$ (6)
where $e_{\mathrm{sem}}\in\mathbb{R}^{d}$ carries the overall semantics. Similarly, we extract content-specific embedding $e_{c}$ and style-specific embedding $e_{s}$ associated with <c> and <s> respectively, enabling isolated semantic guidance for each branch.
Following the standard LoRA parametrization, each layer $i$ is updated by low-rank matrices. To bias the network towards disentanglement, we allocate disjoint index sets $I_{c}$ and $I_{s}$ corresponding to content- and style-sensitive layers: lower and middle layers primarily capture structure and identity (content), while higher layers encode textures and rendering patterns (style). Content-specific matrices $A_{i}^{(c)}$ and style-specific matrices $A_{i}^{(s)}$ are trained with the semantic embedding, and updates are restricted as
$\Delta W_{i}=\begin{cases}B_{i}\,A_{i}^{(c)}(e_{\mathrm{sem}}),&i\in I_{c},\\ B_{i}\,A_{i}^{(s)}(e_{\mathrm{sem}}),&i\in I_{s},\\ 0,&\text{otherwise},\end{cases}$ (7)
so that gradient updates are confined to designated subsets guided by markers, which biases the learned representations towards more independent content and style subspaces.
At inference, the Expert Encoder processes the prompt and produces control scalars $\gamma_{c},\gamma_{s}$ that regulate the strength of content and style adapters. By default, $\gamma_{c},\gamma_{s}\in\{0,1\}$ are set based on marker presence, and users can also continuously adjust them within $[0,1]$ for fine-grained control over content and style intensity. The aggregated backbone at inference is then computed as
$W^{\mathrm{agg}}=W_{\text{init}}+\sum_{i\in I_{c}}\mathsf{E}_{i}\!\left(\gamma_{c}\,\Delta W^{(c)}_{i}\right)+\sum_{i\in I_{s}}\mathsf{E}_{i}\!\left(\gamma_{s}\,\Delta W^{(s)}_{i}\right)$, (8)
where $W_{\text{init}}$ denotes the rank-limited backbone obtained in Eqs.3,4. The operator $\mathsf{E}_{i}(\cdot)$ injects its input matrix into the slot of the $i$-th layer while filling zeros elsewhere, so the summation is a layerwise direct sum rather than a cross-layer addition.
If markers <c> or <s> are absent, the corresponding branch remains inactive. Users may also directly adjust $\gamma_{c},\gamma_{s}$ to control the relative influence of content and style. This enables flexible compositions such as “preserve content but replace style,” without retraining. The full inference procedure is summarized in Algorithm 1, showing how prompts route into disjoint LoRA branches and produce the aggregated backbone $W_{\mathrm{agg}}$.
Figure 4: Visual comparison of content-style combinations. The figure presents a systematic evaluation of different methods in combining specific content elements with distinct artistic styles. Each column showcases the content and style reference followed by the outputs generated. Prompts follow the format “A [content] <c> in [style] <s>”. Competing methods often fail to simultaneously preserve structure and render style, whereas our method produces consistent and coherent content–style compositions.
3.4.2 Asymmetric Classifier-Free Guidance
In standard classifier-free guidance (CFG), the conditional and unconditional paths share the same model weights. When LoRA updates are applied, this causes the unconditional path to be contaminated by content and style adapters, leading to instability in image generation. To address this issue, we propose Asymmetric Classifier-Free Guidance (ACFG), where the conditional path uses LoRA-adapted weights while the unconditional path remains anchored to the rank-limited backbone $W_{\text{init}}$. This asymmetry yields a cleaner and more stable guidance signal without requiring additional training.
Let $i\in\{1,\ldots,L\}$ index model layers and $t\in\{1,\ldots,T\}$ diffusion timesteps. Instead of uniformly applying LoRA updates at all timesteps, we introduce time-dependent activation schedules for content and style: $\gamma_{c}(t)=\mathbb{1}_{t\in T_{c}},\gamma_{s}(t)=\mathbb{1}_{t\in T_{s}}.$ The conditional weights at timestep $t$ are defined as
$W^{\mathrm{cond}}(t)=W_{\text{init}}+\sum_{i\in I_{c}}\mathsf{E}_{i}\!\bigl(\gamma_{c}(t)\,\Delta W_{i}^{(c)}\bigr)+\sum_{i\in I_{s}}\mathsf{E}_{i}\!\bigl(\gamma_{s}(t)\,\Delta W_{i}^{(s)}\bigr),$ (9)
where $W^{\text{init}}_{i}$ denotes the rank-limited backbone, and $\Delta W_{i}^{(c)},\Delta W_{i}^{(s)}$ are content and style LoRA updates. While the unconditional path uses only the initialization backbone:
$W_{i}^{\mathrm{uncond}}(t)=W^{\text{init}}_{i},\quad\forall\,t,i.$ (10)
The conditional prediction is $\epsilon_{\mathrm{cond}}=\epsilon_{\theta}(x_{t}\mid W^{\mathrm{cond}}(t),\mathrm{cond})$, and the unconditional prediction is $\epsilon_{\mathrm{uncond}}=\epsilon_{\theta}(x_{t}\mid W^{\mathrm{uncond}},\varnothing)$. Then the final guided estimate is
$\epsilon_{\theta}^{\mathrm{acfg}}=(1+\omega)\,\epsilon_{\mathrm{cond}}-\omega\,\epsilon_{\mathrm{uncond}},$ (11)
where $\omega$ controls guidance strength. By keeping the unconditional path free of LoRA updates, ACFG avoids interference between content and style adapters and provides a more reliable guidance signal.
Figure 5: Extended visual results. (a) Content–style generations augmented with additional prompt descriptions e.g. “catching a frisbee”, “wearing a hat”, and “driving a car”. The results show that our method flexibly integrates dynamic semantics while preserving the specified content and style. (b) Single-branch generations where only the content LoRA or only the style LoRA is activated. Content-only generations preserve object identity across neutral renderings, while style-only generations transfer artistic features across different subjects, demonstrating disentanglement and controllability.
4 Experiments and Results
4.1 Implementation Details
For content concepts, we use the DreamBooth dataset  [27], containing 30 themes with 4-5 images each. Style concepts come from the StyleDrop  [33] image set, supplemented with classical and contemporary art style images. Importantly, we train each LoRA module (content or style) using only one reference image. To enhance content-style separation during backbone fine-tuning, we construct specialized datasets from our main content and style sources. These datasets emphasize either content preservation or style transfer properties through a controlled generation process based on frequency separation, as content typically resides in lower frequencies and style in higher frequencies. This approach creates clear content-style pairs that provide disentangled signals during backbone fine-tuning, with details provided in the Appendix.Appendix.
All experiments use the Stable Diffusion XL base model  [25] on NVIDIA 4090 GPUs. Our evaluation includes both locally trained LoRAs following our procedure and publicly available pre-trained LoRAs.For locally trained LoRAs, we implement DreamBooth training with rank ($r$) = 64, optimizing weights using Adam-8bit for 1000 steps with batch size 1 and learning rate $1\text{\times}{10}^{-5}$. Training uses minimal prompts like ”A [identifier] photo”, where ”[identifier]” (i.e. ”[c]” for content, ”[s]” for style) serves as a placeholder token representing the specific concept.
(a) Quantitative Comparison
Method
Content Sim.
(CLIP-I) $\uparrow$
Style Sim.
(CLIP-I) $\uparrow$
Combination Score
(GPT-4o) $\uparrow$
Direct Merging 0.65 0.60 0.62
StyleDrop 0.68 0.76 0.70
ZipLoRA 0.70 0.69 0.73
KLoRA 0.71 0.72 0.76
BLoRA 0.74 0.70 0.77
Ours 0.79 0.80 0.83
(b) Ablation study over key components.
Rank-limited Backbone Fine-tuning
Setting Content Sim. Style Sim. Cross-Influence Score $\downarrow$
w/o 0.73 0.69 0.172
w/ 0.82 0.78 0.087
Asymmetric Classifier-Free Guidance
Setting Content Sim. Style Sim. Combination Score $\uparrow$
w/o 0.77 0.64 0.75
w/ 0.84 0.78 0.88
(c) User Study
Method Content Fidelity $\uparrow$ Style Fidelity $\uparrow$ Coherence $\uparrow$
Direct Merging 45.0% 40.0% 42.5%
StyleDrop 55.0% 72.5% 60.0%
ZipLoRA 62.5% 65.0% 67.5%
KLoRA 65.0% 70.0% 70.0%
BLoRA 70.0% 67.5% 72.5%
Ours 82.5% 85.0% 87.5%
Figure 6: Quantitative evaluation of our method. (a) Comparison against baselines using CLIP-I similarities and GPT-4o scores. (b) Ablation study analyzing the effects of Rank-limited Backbone Fine-tuning and Asymmetric CFG. (c) User study results on a 1–5 scale (30 participants, 50 samples), showing clear preference for our method across content fidelity, style fidelity, and overall coherence.
4.2 Results and Analysis
In this section, we present a comprehensive analysis of our proposed method, encompassing visual demonstrations, quantitative evaluations against baselines, and an ablation study investigating the contributions of key components.
Visual Results. Figure 4 compares our method with representative LoRA fusion baselines, including ZipLoRA, BLoRA, and KLoRA. Prompts are formatted as “A [content] <c> in [style] <s>” with different content–style pairs such as dogs, plush toys, and objects. While existing methods often compromise either content fidelity e.g. structural distortion or style rendering e.g. muted patterns, our approach produces outputs that preserve object identity while faithfully capturing artistic styles. Figure 5 provides extended qualitative demonstrations. The left part shows content–style combinations augmented with additional prompt descriptions e.g. “catching a frisbee”, “playing with a ball”, or “driving a car”. Our method integrates these dynamic semantics while faithfully maintaining the designated content and style. The right part isolates the effect of single-branch adapters. Activating only the content branch preserves structural fidelity of objects across neutral renderings, while activating only the style branch applies stylistic transformations consistently across varied subjects. These observations confirm that our framework achieves effective disentanglement and controllability during inference.
Quantitative Evaluation. We evaluate our method using both automatic and human-centered metrics. For automatic evaluation, we adopt Content Similarity and Style Similarity based on CLIP-I embeddings, as well as a Combination Score derived from GPT-4o that measures the overall coherence of content–style integration. Each score is averaged over 10 distinct content–style pairs, with 10 generated images per pair. As shown in Table 6(a), baseline methods exhibit different trade-offs: Direct Merging fails to preserve either content or style, StyleDrop emphasizes style at the expense of structure, and ZipLoRA, KLoRA, and BLoRA achieve more balanced results but remain limited in overall coherence. Our method consistently achieves the highest scores across all metrics, demonstrating its ability to preserve structural fidelity while maintaining faithful style transfer and coherent integration.
Ablation and User Study. Table 6(b) analyzes the contributions of our two key components. Rank-limited backbone fine-tuning improves CLIP-I content/style similarity and reduces cross-influence, indicating stronger content–style disentanglement. The Cross-Influence score measures the relative drop in content/style similarity when composing both factors versus single-factor baselines. Specifically, it is computed as the product of CLIP-I similarities between the generated image and both reference images; lower values indicate better disentanglement with reduced mutual interference. Asymmetric CFG further strengthens integration quality, reflected by higher human-aligned combination scores and consistent gains in CLIP-I. Table 6(c) further reports results from a user study with 30 participants. Participants consistently preferred our method across content fidelity, style fidelity, and overall coherence, aligning with the automatic metrics. This combination of quantitative and perceptual evidence confirms that our framework achieves superior disentanglement, fidelity, and integration compared with prior approaches.
5 Conclusion
We presented a framework for personalized image generation that integrates efficient adaptation strategies. Rank-constrained backbone fine-tuning encourages distinct subspaces, providing a inductive bias for disentanglement. An expert encoder with selective aggregation enhances semantic guidance, enabling flexible composition, and timestep-dependent asymmetric CFG introduces a training-free sampling strategy that improves visual quality.
Our experiments substantiate these contributions, demonstrating a measurable reduction in content–style cross-influence, consistent improvements across automatic evaluation metrics, and clear user preference in perceptual studies. Limitations include architecture-dependent layer selection, reliance on text embedding quality for semantic extension, and minor overhead from timestep-dependent correction. Future work will explore automated layer assignment and extensions to broader diffusion models and modalities.
Supplementary Material
Appendix
Appendix A Trunk (Backbone) Fine-tuning Details
A.1 Targeted Dataset Construction
We construct a dataset of 100 content–style contrast pairs, $\mathcal{D}=\{(I_{c}^{k},I_{s}^{k})\}_{k=1}^{100}$. This dataset is used to fine-tune the backbone, encouraging content–style disentanglement and providing the base weights ($W_{\text{init}}$) for subsequent LoRA adapter training. The pairs are generated using a pre-trained diffusion model with frequency-based controls.
For each pair $(I_{c}^{k},I_{s}^{k})\in\mathcal{D}$: image $I_{c}^{k}$ emphasizes content, guided by a content prompt $P_{c}^{k}$ and a style modifier $P_{sm}^{k}$; image $I_{s}^{k}$ emphasizes style, guided by a style prompt $P_{s}^{k}$ and a content modifier $P_{cm}^{k}$. One specific $P_{sm}^{k}$ and one $P_{cm}^{k}$ are chosen per pair to form a one-to-one contrast.
We employ frequency-domain manipulations during iterative denoising. Let $\mathcal{F}$ denote a frequency transform (e.g., Discrete Cosine Transform) and $\mathcal{F}^{-1}$ its inverse. Masks $\mathcal{M}_{\text{low}}$ and $\mathcal{M}_{\text{high}}$ preserve low- and high-frequency components, respectively. With noisy latent $z_{t}$ at timestep $t$ and text embedding $e=\mathrm{TextualEmbeddings}(P)$, the process is:
$\hat{z}_{0}$ $=\mathrm{Predictor}(z_{t},t,e),\quad\text{(Predict clean latent)}$
$z_{t-1}^{\text{filtered}}$ $=\mathcal{F}^{-1}\!\big(\mathcal{M}_{\text{low}}\odot\mathcal{F}(\hat{z}_{0})\big)\quad\text{(for }I_{c}^{k}\text{ generation)},$
$z_{t-1}^{\text{filtered}}$ $=\mathcal{F}^{-1}\!\big(\mathcal{M}_{\text{high}}\odot\mathcal{F}(\hat{z}_{0})\big)\quad\text{(for }I_{s}^{k}\text{ generation)},$
$z_{t-1}$ $=\mathrm{UpdateRule}(z_{t},z_{t-1}^{\text{filtered}},\dots),$
where $\odot$ denotes element-wise multiplication. This design concentrates variation along the intended factor (content or style).
Table 1: Conceptual illustration of content–style contrast pair generation. For each pair $(I_{c}^{k},I_{s}^{k})$, one varying modifier is chosen for $I_{c}^{k}$ and one for $I_{s}^{k}$.
Property $I_{c}^{k}$ (content fixed) $I_{s}^{k}$ (style fixed)
Target image Generated with fixed content Generated with fixed style
Base prompt $P_{c}^{k}$: “a red car” $P_{s}^{k}$: “in the style of Van Gogh”
Varying modifier $P_{sm}^{k}\in\{\text{watercolor},\ \text{oil painting},\dots\}$ $P_{cm}^{k}\in\{\text{starry night},\ \text{sunflower},\dots\}$
Dominant frequency filter Low $\mathcal{M}_{\text{low}}$ High $\mathcal{M}_{\text{high}}$
Resulting characteristic Content preserved with varied textures Style preserved with varied subjects
A.2 Training Settings
We fine-tune the backbone weights $\{W_{l}\}_{l=1}^{L}$ by minimizing:
$\mathcal{L}_{\text{trunk}}=\mathcal{L}_{\text{task}}(\{W_{l}\},\mathcal{D})+\lambda_{\text{reg}}\sum_{l=1}^{L}\|B_{l}\|_{F}^{2},$ (12)
where $B_{l}$ is a learnable basis for each layer $l$. Let $B_{l}=Q_{l}R_{l}$ be the QR decomposition of the basis, where $Q_{l}$ contains the orthonormal vectors ($Q_{l}^{\top}Q_{l}=I$). Consistent with the update rule discussed in the main text, the weights are updated as:
$W_{l}=W_{l}^{(0)}-Q_{l}Q_{l}^{\top}W_{l}^{(0)}.$ (13)
This update effectively projects the original weights $W_{l}^{(0)}$ onto the orthogonal complement of the learned subspace $\mathrm{span}(Q_{l})$. We denote the resulting post-finetuning backbone as $W_{\text{init}}$ (also written $\widetilde{W}$ in figures). The update $\Delta W_{l}=-Q_{l}Q_{l}^{\top}W_{l}^{(0)}$ lies in $\mathrm{span}(Q_{l})$.
The per-layer rank $r_{l}$ for the basis $B_{l}$ follows a linear schedule:
$r_{l}=r_{\max}-\frac{l-1}{L-1}\big(r_{\max}-r_{\min}\big),\quad l=1,\dots,L,$ (14)
with maximum rank $r_{\max}=128$ and minimum rank $r_{\min}=4$.
$\mathcal{L}_{\text{task}}(\{W_{l}\},\mathcal{D})=\frac{1}{N}\sum_{k=1}^{N}\mathbb{E}_{t,z_{t}}\Bigg[\big\|G(\{W_{l}\},z_{t},t,\mathrm{TextualEmbeddings}(P_{c}^{k},P_{sm}^{k}))-I_{c}^{k}\big\|_{1}\\ +\big\|G(\{W_{l}\},z_{t},t,\mathrm{TextualEmbeddings}(P_{cm}^{k},P_{s}^{k}))-I_{s}^{k}\big\|_{1}\Bigg]+\alpha\,\mathcal{L}_{\text{perceptual}}(\{W_{l}\},\mathcal{D}).$ (15)
$\mathcal{L}_{\text{perceptual}}=\frac{1}{N}\sum_{k=1}^{N}\mathbb{E}_{t,z_{t}}\sum_{j\in\mathrm{VGG\_layers}}\frac{1}{M_{j}}\Bigg[\big\|\phi_{j}\big(G(\{W_{l}\},z_{t},t,\mathrm{TextualEmbeddings}(P_{c}^{k},P_{sm}^{k}))\big)-\phi_{j}(I_{c}^{k})\big\|_{1}\\ +\big\|\phi_{j}\big(G(\{W_{l}\},z_{t},t,\mathrm{TextualEmbeddings}(P_{cm}^{k},P_{s}^{k}))\big)-\phi_{j}(I_{s}^{k})\big\|_{1}\Bigg].$ (16)
The task loss $\mathcal{L}_{\text{task}}$ (Eq. 15) is computed over the dataset $\mathcal{D}$. Here, $N=100$, $G(\cdot)$ represents the $\hat{x}_{0}$ predictor (i.e., the U-Net), $t$ is a sampled timestep, and $z_{t}$ is the corresponding noisy latent created from the target image.
The perceptual term $\mathcal{L}_{\text{perceptual}}$ (Eq. 16) is defined as: where $\phi_{j}$ are features from VGG-19 (relu1_1 to relu5_1) and $M_{j}$ is the number of elements in the feature map. We set $\alpha=0.1$ and $\lambda_{\text{reg}}=$1\text{\times}{10}^{-4}$\$.
Optimization uses the AdamW optimizer with a cosine decay learning rate schedule and a 500-step warm-up. The learning rate starts at $1\text{\times}{10}^{-5}$, warms up from $1\text{\times}{10}^{-6}$, and decays to a minimum of $1\text{\times}{10}^{-7}$. We use a batch size of 8 and train for 5000 steps.
Table 2: Summary of backbone fine-tuning hyperparameters.
Parameter Value
Optimizer AdamW
Initial learning rate $1\text{\times}{10}^{-5}$
LR schedule Cosine decay with 500 warm-up steps
Minimum learning rate $1\text{\times}{10}^{-7}$
Batch size 8
Optimization steps 5000
Tuned U-Net layers Attention and standard convolutional layers
Rank schedule $r_{l}$ (Eq. 14) Linear; $r_{\max}=128$, $r_{\min}=4$
$\mathcal{L}_{\text{task}}$ components $\ell_{1}$ + Perceptual (VGG-19 relu1_1 to relu5_1)
Perceptual loss weight $\alpha$ 0.1
$B_{l}$ regularization $\lambda_{\text{reg}}$ $1\text{\times}{10}^{-4}$
Hardware 2 NVIDIA 4090 GPUs
Appendix B LoRA Training and Inference Details
B.1 Setting and Notation
LoRA adapters are trained on the U-Net backbone initialized with the pre-tuned weights $W_{\text{init}}$ (from Eq. 13). We define disjoint sets of U-Net layers $I_{c}$ (for content) and $I_{s}$ (for style), satisfying $I_{c}\cap I_{s}=\emptyset$. The adapters are trained only on the layers specified in $I_{c}\cup I_{s}$.
LoRA parameterization. For any layer $i\in I_{c}\cup I_{s}$, we use a fixed LoRA rank $r=16$. Content adapters learn parameters $\{B_{i}^{(c)},A_{i}^{(c)}\}$ for $i\in I_{c}$; style adapters learn parameters $\{B_{i}^{(s)},A_{i}^{(s)}\}$ for $i\in I_{s}$. The update is $\Delta W_{i}=B_{i}A_{i}$.
Training procedure. During content adapter training, gradients for $\{B_{i}^{(c)},A_{i}^{(c)}\}$ are enabled only for $i\in I_{c}$ (and masked for $i\in I_{s}$). Conversely, style adapter training only updates parameters for $i\in I_{s}$. Training hyperparameters are listed in Table 3. The specific layer assignments for content and style are shown in Table 4.
Table 3: LoRA adapter training hyperparameters.
Parameter Value
Optimizer AdamW
Learning rate $1\text{\times}{10}^{-5}$
Training steps 1000–2000
Batch size 1
Backbone host $W_{\text{init}}$ (from Eq. 13)
Table 4: U-Net layer assignments for content $(I_{c})$ and style $(I_{s})$ adapters.
Content layers $(I_{c})$ Style layers $(I_{s})$
down_blocks.2.attentions.0; mid_block.attentions.0; up_blocks.0.attentions.1; up_blocks.1.attentions.0; up_blocks.2.attentions.0 down_blocks.0.attentions.0; down_blocks.1.resnets.0; mid_block.resnets.1; down_blocks.0.resnets.1; down_blocks.1.attentions.1
Expert encoder and aggregation. Three MLP encoders, $E_{t},E_{c},E_{s}$, produce 64-dimensional embeddings from a concept ID embedding, a CLIP content text embedding, and a CLIP style text embedding, respectively. Their concatenated features are fed into an aggregation head that outputs nonnegative scaling factors $(\gamma_{c},\gamma_{s})$. At inference, the aggregated weights are:
$W^{\mathrm{agg}}=W_{\text{init}}+\sum_{i\in I_{c}}\mathsf{E}_{i}\!\big(\gamma_{c}\,\Delta W^{(c)}_{i}\big)+\sum_{i\in I_{s}}\mathsf{E}_{i}\!\big(\gamma_{s}\,\Delta W^{(s)}_{i}\big)$,
where $\mathsf{E}_{i}(\cdot)$ is an operator that injects the low-rank update $\Delta W_{i}=B_{i}A_{i}$ at layer $i$.
B.2 Disentanglement Metrics and Results
Table 5: Disentanglement evaluation metrics. $\mathcal{G}(p_{c},p_{s})$ denotes the final image generated from content prompt $p_{c}$ and style prompt $p_{s}$. Higher is better for $\mathcal{S}_{c},\mathcal{S}_{s}$; lower is better for $\mathcal{S}_{x}$.
Metric Formula Description
Content preservation $\mathcal{S}_{c}$ $\frac{1}{N_{s}}\sum_{j=1}^{N_{s}}\mathrm{sim}_{c}(\mathcal{G}(p^{\prime}_{c},p_{s}^{j}),p^{\prime}_{c})$ Similarity to target content
Style fidelity $\mathcal{S}_{s}$ $\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\mathrm{sim}_{s}(\mathcal{G}(p_{c}^{i},p^{\prime}_{s}),p^{\prime}_{s})$ Similarity to target style
Cross-influence $\mathcal{S}_{x}$ $\frac{1}{N_{c}N_{s}}\sum_{i,j}\mathrm{interference}(p_{c}^{i},p_{s}^{j})$ Undesired cross effect
Table 6: Quantitative comparison of content–style disentanglement.
Method $\mathcal{S}_{c}\uparrow$ $\mathcal{S}_{s}\uparrow$ $\mathcal{S}_{x}\downarrow$
Standard LoRA 0.75 0.70 0.42
Trunk fine-tune + layer-selective LoRA (Ours) 0.90 0.88 0.18
B.3 Semantic Extension Examples
Table 7: Types of semantic extension (generalization).
Type Condition Description
I $p^{\prime}_{c}\neq p_{c}^{i},\ p^{\prime}_{s}=p_{s}^{j}$ New content, trained style
II $p^{\prime}_{c}=p_{c}^{i},\ p^{\prime}_{s}\neq p_{s}^{j}$ Trained content, new style
III $p^{\prime}_{c}\neq p_{c}^{i},\ p^{\prime}_{s}\neq p_{s}^{j}$ New content, new style
Table 8: Semantic extension examples via prompt variation.
Trained content Trained style Inference prompts Type
“a red car” “Van Gogh style” $p^{\prime}_{c}$: “a blue car”, $p^{\prime}_{s}$: “Van Gogh style” I
“a red car” “Van Gogh style” $p^{\prime}_{c}$: “a red car”, $p^{\prime}_{s}$: “sketch style” II
“a red car” “Van Gogh style” $p^{\prime}_{c}$: “a bicycle”, $p^{\prime}_{s}$: “sketch style” III
“person ID X” “pixel art” $p^{\prime}_{c}$: “person ID Y”, $p^{\prime}_{s}$: “pixel art” I
“person ID X” “pixel art” $p^{\prime}_{c}$: “person ID X”, $p^{\prime}_{s}$: “cyberpunk style” II
Figure 7: Visual results demonstrating content and style adapter performance.
Appendix C Asymmetric Classifier-Free Guidance (ACFG) for Timestep-Dependent Aggregation
We introduce Asymmetric Classifier-Free Guidance (ACFG) during the sampling process. In ACFG, the conditional branch (guidance-seeking) uses the dynamically aggregated LoRA weights, while the unconditional branch (guidance-free) always uses the static, pre-tuned backbone initialization $W_{\text{init}}$.
Conditional weights.
For timestep $t\in\{1,\dots,T\}$, the weights for layer $i$ are:
$W_{i}^{\mathrm{cond}}(t)=W_{\text{init},i}+\gamma_{c}(t)\,\Delta W^{(c)}_{i}+\gamma_{s}(t)\,\Delta W^{(s)}_{i}$, (17)
where $\gamma_{c}(t),\gamma_{s}(t)\geq 0$ are dynamic schedules (which can be binary or continuous). The unconditional weights are fixed:
$W_{i}^{\mathrm{uncond}}(t)=W_{\text{init},i}\qquad\forall i,t.$ (18)
Guided prediction.
Let $\epsilon_{\mathrm{cond}}=\epsilon_{\theta}(x_{t}\mid\{W_{i}^{\mathrm{cond}}(t)\},\mathrm{cond})$ be the noise estimate using conditional weights and $\epsilon_{\mathrm{uncond}}=\epsilon_{\theta}(x_{t}\mid\{W_{i}^{\mathrm{uncond}}(t)\},\varnothing)$ be the estimate using unconditional weights. The final ACFG prediction is:
$\epsilon_{\theta}^{\mathrm{acfg}}(t)=(1+\omega)\,\epsilon_{\mathrm{cond}}-\omega\,\epsilon_{\mathrm{uncond}}.$ (19)
where $\omega$ is the guidance scale.
Figure 8: Additional visual results, including ACFG comparisons and semantic extensions.
Zero-shot temporal scaling.
We propose a smooth schedule
$\alpha(t)=\alpha_{\min}+(\alpha_{\max}-\alpha_{\min})\,g\!\left(\frac{T-t}{T}\right),$
with a monotone function $g:[0,1]\to[0,1]$ (e.g., a cosine schedule). This $\alpha(t)$ modulates the aggregation weights $\gamma_{c}(t)$ and $\gamma_{s}(t)$ to emphasize content early in the sampling (high $t$) and refine style later (low $t$), while always keeping the unconditional branch fixed as in Eq. 18.
Table 9: Performance by content–style pairing with ACFG. Metric shown is a composite quality score (higher is better).
Content–style pairing Fixed-weights baseline ACFG (ours) Relative improvement
Similar semantics 0.81 0.87 +7.4%
Distant semantics 0.68 0.79 +16.2%
Complex composition 0.65 0.80 +23.1%
Table 10: Ablation of ACFG components. Higher indicates better quality.
Component configuration Quality score Relative change
Full ACFG as proposed 0.86 Reference
Constant temporal scaling (no time dependence) 0.78 $9.3\%$
No low-rank constraint on temporal adjustments 0.82 $4.7\%$
Standard CFG (uncond. path uses cond. weights) 0.75 $12.8\%$
Appendix D Additional Visualizations
Figures 7 and 8 provide additional qualitative results, complementing the visualizations in the main paper. These examples cover diverse objects, complex scenes, and challenging artistic styles, further demonstrating the robustness of our method. We also include failure cases to illustrate current limitations, such as difficulty with highly abstract styles or extreme compositional changes.

### v2

@@
 1 Introduction
-Recent advancements in deep learning have fueled remarkable progress in generative models, particularly within the domain of personalized image generation [11, 2, 30].
-Unlike generic image synthesis, personalized generation focuses on customizing outputs based on user-provided references, such as a specific subject’s appearance or stylistic attributes.
-In more sophisticated scenarios, these systems aim to simultaneously preserve subject identity while emulating a target style, enabling applications in creative design, entertainment, digital avatars, personalized marketing, and beyond.
-Among the various techniques developed for personalized image generation, Low-Rank Adaptation (LoRA) has emerged as a particularly prominent and efficient approach [15].
-LoRA facilitates adaptation by introducing small, trainable low-rank matrices into specific layers of large pre-trained text-to-image diffusion models, thereby allowing for efficient fine-tuning on new concepts using only a small number of references.
-When integrated with personalization frameworks like DreamBooth, LoRA-based methods enable the customization of generated images to specific content, e.g.
-preserving the identity of a person, or style, e.g.
-emulating a particular artistic rendering or texture, making them widely adopted [29].
-Moreover, LoRA exhibits a unique compositional property: multiple independently trained LoRA modules, each corresponding to different visual concepts, can be simultaneously applied during inference.
-This enables the model to generate images that blend diverse attributes, such as combining a specific subject with a distinct artistic style.
-However, the combinatorial use of LoRA weights introduces new challenges.
-Naively merging multiple LoRA modules often leads to degraded visual quality or semantic inconsistency, due to entangled or conflicting latent representations.
-Consequently, recent works have explored various strategies for more effective LoRA fusion, including approaches that re-optimize weights for specific combinations [31], adjust internal attention mechanisms [9], or dynamically modulate different LoRAs’ influence during generation [22].
-Despite recent progress, existing LoRA combination strategies still face several fundamental challenges, which motivate the improvements proposed in this paper.
+Recent advancements in deep learning have fueled remarkable progress in generative models, particularly within personalized image generation [11, 2, 33].
+Unlike generic image synthesis, personalized generation focuses on customizing outputs based on user-provided references, such as a specific subject’s appearance or stylistic attributes, enabling applications in creative design, entertainment, digital avatars, and personalized marketing.
+Among the various techniques developed for personalized image generation, Low-Rank Adaptation (LoRA) [15] has emerged as a particularly efficient approach by introducing small, trainable low-rank matrices into pre-trained text-to-image diffusion models, allowing fine-tuning on new concepts using only a small number of references.
+When integrated with frameworks like DreamBooth, LoRA-based methods enable customization to specific content or style from minimal examples [32].
+Moreover, multiple independently trained LoRA modules can be simultaneously applied during inference, enabling images that blend diverse attributes such as combining a specific subject with a distinct artistic style.
+However, naively merging multiple LoRA modules often leads to degraded quality or semantic inconsistency due to entangled representations, motivating recent work on more effective fusion strategies [34, 9, 25].
+Despite recent progress, existing LoRA combination strategies still face several fundamental challenges.
 First, although pre-trained diffusion models empirically exhibit some degree of composability with respect to LoRA modules, these models are not explicitly trained to support such combinations.
-As a result, directly merging LoRA weights often fails to achieve a clean disentanglement between content and style factors present in the reference images [19, 18, 1].
-Inspired by the meta-learning paradigm, particularly Model-Agnostic Meta-Learning (MAML) [8, 6], which seeks optimal initializations for fast and generalizable task adaptation, we propose a novel network re-initialization strategy for pre-trained diffusion models that enhances their compatibility with compositional LoRA usage.
-Specifically, we introduce a low-rank constrained adaptation phase over the parameters of the pre-trained diffusion model inspired by PaRa [5].
-The additional training stage narrows the model’s generative space, shifting it from a broadly generalized prior to a more compact and disentangled representation space that better supports the fusion of multiple LoRA modules.
-Another critical limitation lies in the inherent granularity ambiguity associated with the definition of a “subject” in personalized image generation.
-In practical scenarios, the notion of a subject often encompasses a hierarchy of visual attributes—for instance, a portrait image may include multiple levels of subject-specific features, e.g.
+As a result, directly merging LoRA weights often fails to achieve clean disentanglement between content and style factors [21, 20, 1].
+Inspired by the meta-learning paradigm, particularly MAML [8, 6], which seeks optimal initializations for fast and generalizable task adaptation, we introduce a low-rank constrained adaptation phase inspired by PaRa [5] that narrows the model’s generative space to a more compact and disentangled representation that better supports multi-LoRA fusion.
+Second, the notion of a subject in personalized image generation often encompasses a hierarchy of visual attributes, e.g.
 facial identity, hairstyle, clothing style, and accessory details.
-However, current methods typically collapse this rich semantic structure into a single coarse-grained token representation within the input prompt [24, 20].
-This oversimplification neglects the internal structure of subject attributes and lacks the mechanisms to differentiate or control the preservation and variation of fine-grained elements.
-To address this, we propose a novel expert encoder system that enables fine-grained control over subject features at different levels.
-Our system implements specially designed content and style adapters operating in disjoint network layers.
-The expert encoder processes identity labels, content descriptions, and style specifications through separate branches, facilitating selective weights aggregation based on semantic relevance The designed expert system enables precise delineation of feature preservation hierarchies through prompts, allowing for discriminative activation of specific subject attributes at variable granularity levels according to specified semantic constraints.
-Finally, existing LoRA fusion strategies, such as ZipLoRA [31], typically rely on additional optimization procedures to reconcile differences between LoRA modules [37, 12].
-While effective to some extent, directly modifying LoRA parameters during fusion can inadvertently alter or suppress critical elements encoded in the original weights, leading to loss of identity or stylistic fidelity.
-Moreover, these methods incur non-trivial computational overhead.
-To overcome these limitations, we propose a novel tuning-free LoRA combination method that circumvents the need for parameter-level fusion.
-At the core of our approach is our developed asymmetric Classifier-Free Guidance (CFG) mechanism applied during the diffusion process.
-Rather than adjusting LoRA weights directly, we selectively apply time-dependent modifications only to the conditional path where combined LoRA weights are applied, while keeping the unconditional path that uses only the base model without LoRA adapters fixed to the base model weights.
-Our asymmetric design yields a dynamic guidance signal that effectively steers the generation process toward a desirable balance between content and style, without introducing additional training cost or risking the degradation of essential features, thereby offering a lightweight and robust alternative to existing LoRA fusion techniques.
-As illustrated in Figure 2, our proposed unified framework integrates rank-constrained fine-tuning to learn decoupled content and style subspaces, leverages prompt-guided mechanisms for semantic control and extension via selective adapter aggregation guided by an expert encoder, and employs a time-step-aware guidance correction scheme for stable, high-fidelity generation without extra training.
+However, current methods typically collapse this rich semantic structure into a single coarse-grained token representation [27, 22], neglecting the internal structure of subject attributes and lacking mechanisms to control the preservation of fine-grained elements.
+To address this, we propose an expert encoder system with specially designed content and style adapters operating in disjoint network layers.
+The expert encoder processes identity labels, content descriptions, and style specifications through separate branches, facilitating selective aggregation based on semantic relevance and enabling fine-grained control over feature preservation at different levels.
+Third, existing LoRA fusion strategies such as ZipLoRA [34] typically rely on additional optimization procedures to reconcile differences between LoRA modules [41, 12].
+Directly modifying LoRA parameters during fusion can inadvertently alter or suppress critical elements encoded in the original weights, leading to loss of identity or stylistic fidelity, while also incurring non-trivial computational overhead.
+To overcome these limitations, we propose a tuning-free asymmetric Classifier-Free Guidance (CFG) mechanism.
+Rather than adjusting LoRA weights directly, we selectively apply time-dependent modifications only to the conditional path while keeping the unconditional path anchored to base model weights, yielding a dynamic guidance signal that steers generation toward a desirable content-style balance without additional training cost.
+As illustrated in Figure 2, our proposed unified framework integrates rank-constrained fine-tuning to learn decoupled content and style subspaces, leverages prompt-guided mechanisms for semantic control via selective adapter aggregation guided by an expert encoder, and employs a timestep-aware guidance correction scheme for stable, high-fidelity generation without extra training.
 To summarize, we make the following key contributions:
@@
 Consequently, current stylization pipelines often struggle to integrate diverse visual concepts without compromising either semantic coherence or stylistic integrity  [39, 26].
-2.2 Stylization through Fine-Tuning Methods
+2.2 Fine-Tuning and LoRA Composition for Stylization
 The rapid advancement of text-to-image diffusion models has spurred diverse personalization strategies.
-Full-model fine-tuning approaches, exemplified by DreamBooth [27], can achieve high-fidelity style transfer but incur substantial computational costs and require lengthy retraining on domain-specific examples.
-To mitigate these demands, Parameter-Efficient Fine-Tuning techniques have emerged [28].
-Among these, LoRA introduces small and trainable low-rank decomposition matrices, injecting new stylistic or semantic concepts while keeping the majority of the pre-trained model weights frozen, thus striking a balance between adaptability and efficiency.
-Complementing weight-space adaptation methods, token-based optimization techniques, such as Textual Inversion [10], learn new embeddings to represent novel concepts within the model’s token vocabulary.
-Adapter-based schemes offer an alternative pathway, for example, StyleDrop [33] appends lightweight layers to attention blocks, isolating style information in dedicated modules.
-Meta-learning, as a cross-task method for domain transfer, is similar in principle to style transfer in personalized image generation, providing new inspirations for fine-tuning methods [39].
-2.3 Content-Style LoRAs Merging
-Methods for combining multiple LoRA modules, effective for capturing distinct concepts, are explored to achieve simultaneous content and style adaptation.
-Parameter interpolation is a basic approach, which is prone to suboptimal outcomes, especially when parameters conflict [34, 36].
-To address these challenges, several studies focus on refining the composition mechanism or the application process itself.
-ZipLoRA [31] learns column-wise mixing coefficients to better balance style and content elements, although it introduces the need for additional optimization procedures for each new content-style pairing.
-B-LoRA [9], in contrast, differentiates the functional roles within attention modules, aiming for a more effective separation of object and style representations.
-LoRA Composition [40] employs a cyclic update mechanism, allowing LoRA weights to iteratively guide the generation without explicit parameter merging.
+Full-model fine-tuning approaches such as DreamBooth [30] achieve high-fidelity transfer but incur substantial computational costs and require lengthy retraining on domain-specific examples.
+To mitigate these demands, Parameter-Efficient Fine-Tuning (PEFT) techniques have emerged [31, 12].
+Among these, LoRA [15] introduces small, trainable low-rank decomposition matrices, injecting new stylistic or semantic concepts while keeping the majority of pre-trained weights frozen, thus striking a balance between adaptability and efficiency.
+Complementary approaches include token-based optimization techniques such as Textual Inversion [10], which learns new embeddings to represent novel concepts within the model’s token vocabulary, and adapter-based schemes such as StyleDrop [37], which appends lightweight layers to attention blocks to isolate style information in dedicated modules.
+Meta-learning, as a cross-task method for domain transfer, shares similar principles with style transfer in personalized image generation, providing new inspirations for fine-tuning methods [45].
+A key challenge arises when combining multiple LoRA modules for simultaneous content and style adaptation.
+Parameter interpolation [38, 40] provides a basic approach but is prone to suboptimal outcomes, especially when parameters conflict.
+To address this, several studies refine the composition mechanism: ZipLoRA [34] learns column-wise mixing coefficients for each content-style pairing, B-LoRA [9] differentiates functional roles within attention modules for content-style separation, and LoRA Composition [46] employs a cyclic update mechanism to iteratively guide generation without explicit parameter merging.
+Training-free approaches have also been explored, with K-LoRA [25] selectively applying LoRAs to different attention layers and CLoRA [24] proposing contrastive test-time composition.
+Other directions include QR-LoRA [43], which applies QR decomposition for parameter-efficient training, and LoRA.rar [35], which learns merging via hypernetworks.
+While these methods improve the merging or composition process, the upstream entanglement inherent in standard backbones remains underexplored.
 3 Methodology
 3.1 Preliminaries
-LoRA is an efficient fine-tuning method that adapts pretrained diffusion models by adding trainable low-rank matrices $A\in\mathbb{R}^{r\times n}$ and $B\in\mathbb{R}^{m\times r}$ to frozen weights $W_{0}\in\mathbb{R}^{m\times n}$, yielding $W=W_{0}+BA$ with $r\ll\min(m,n)$.
-In personalized image generation, content LoRAs $\Delta W_{c}$ represent subject-specific features like identity and structure, while style LoRAs $\Delta W_{s}$ capture artistic aspects such as texture and color.
-Both can be trained efficiently from a single reference image, and during inference they can be composed into the backbone to form personalized generative elements.
+LoRA is an efficient fine-tuning method that adapts pre-trained diffusion models by adding trainable low-rank matrices $A\in\mathbb{R}^{r\times n}$ and $B\in\mathbb{R}^{m\times r}$ to frozen weights $W_{0}\in\mathbb{R}^{m\times n}$, yielding $W=W_{0}+BA$ with $r\ll\min(m,n)$.
+In personalized image generation, content LoRAs $\Delta W_{c}$ represent subject-specific features such as identity and structure, while style LoRAs $\Delta W_{s}$ capture artistic aspects such as texture and color.
+Both can be trained efficiently from a single reference image, and during inference they can be composed into the backbone to produce personalized outputs.
 3.2 Framework Overview
-We propose CRAFT-LoRA, a unified framework for personalized image synthesis that addresses the challenges of content-style disentanglement and fusion in LoRA-based methods.
+We propose CRAFT-LoRA, a unified framework for personalized image synthesis that addresses content-style disentanglement and fusion in LoRA-based methods.
 As shown in Fig.
-2, the framework includes three main stages: (1) Training Stage, where content and style LoRA weights are learned separately based on a rank-constrained fine-tuned backbone; (2) Prompt Text Guidance, which employs an expert encoder system with specialized branches to provide semantic control over content and style during training and inference; (3) Inference Stage, utilizing a timestep-aware asymmetric classifier-free guidance scheme for stable and flexible fusion of LoRA weights without extra training.
+2, the framework consists of three main stages: (1) Training Stage, where content and style LoRA weights are learned separately based on a rank-constrained fine-tuned backbone; (2) Prompt Text Guidance, which employs an expert encoder system with specialized branches to provide semantic control over content and style during training and inference; (3) Inference Stage, utilizing a timestep-aware asymmetric classifier-free guidance scheme for stable and flexible fusion of LoRA weights without extra training.
 3.3 Rank-Limited Fine-Tuning of Backbone
-As mentioned in the introduction, pre-trained diffusion models do not explicitly support clear disentanglement of content and style factors when adapting weights from a reference image, thus separation enhancement requires additional fine-tuning of the diffusion backbone.
-Motivated by the adaptability of MAML [8, 17] under few-shot constraints and leveraging the rich hierarchical features of SDXL [25], we propose to fine-tune the U-Net backbone under a rank-constrained mechanism.
-Our approach provides optimized initialization and structurally improves the disentanglement of content and style in the U-Net architecture, providing a more effective foundation for subsequent LoRA separation and personalized synthesis.
-3.3.1 Rank Reduction Fine-Tuning.
-Inspired by PaRa method [5], we denote the frozen backbone weight matrix of layer $l$ as $W_{l}^{(0)}\in\mathbb{R}^{d_{\mathrm{out}}^{l}\times d_{\mathrm{in}}^{l}}$, with $L$ layers in total.
+As discussed in the introduction, pre-trained diffusion models do not explicitly support clear disentanglement of content and style factors when adapting weights from a reference image.
+Motivated by the adaptability of MAML [8, 18] under few-shot constraints and leveraging the rich hierarchical features of SDXL [28], we propose to fine-tune the U-Net backbone under a rank-constrained mechanism that provides optimized initialization and structurally improves content-style disentanglement for subsequent LoRA training.
+Rank Reduction Fine-Tuning.
+Inspired by PaRa [5], we denote the frozen backbone weight matrix of layer $l$ as $W_{l}^{(0)}\in\mathbb{R}^{d_{\mathrm{out}}^{l}\times d_{\mathrm{in}}^{l}}$, with $L$ layers in total.
 We introduce a learnable basis matrix $B_{l}\in\mathbb{R}^{d_{\mathrm{in}}^{l}\times r_{l}}$ per layer, where $r_{l}$ is the rank constraint satisfying $r_{l}\ll\min(d_{\mathrm{out}}^{l},d_{\mathrm{in}}^{l})$.
@@
 $W_{l}=W_{l}^{(0)}-Q_{l}Q_{l}^{\top}W_{l}^{(0)}.$ (1)
-This constrains updates to directions orthogonal to the learned low-rank subspace, introducing an architectural bias that encourages reduced overlap between content and style representations.
-We show empirically in Appendix.
-Appendix that it reduces cross-influence.
+This operation constrains updates to directions orthogonal to the learned low-rank subspace, introducing an architectural bias that encourages reduced overlap between content and style representations.
+We show empirically in Appendix Appendix that it reduces cross-influence.
 To further allocate adaptation capacity hierarchically, we schedule:
@@
 assigning higher ranks to earlier layers, which typically encode low-level structural and textural information where content and style factors are more intertwined, thus requiring more adaptation capacity.
+In implementation, we set $r_{\max}=128$ and $r_{\min}=4$.
+Compared to a uniform rank $r=64$ across all layers, this schedule improves Content Similarity by 4.8%, Style Similarity by 5.2%, and reduces cross-influence by 11.3%, confirming the benefit of hierarchical rank allocation.
 For content-style disentanglement, we train separate basis matrices $B_{l}^{\mathrm{content}}$ and $B_{l}^{\mathrm{style}}$, obtaining corresponding orthonormal subspaces $Q_{l}^{\mathrm{content}}$ and $Q_{l}^{\mathrm{style}}$.
-Merging these via concatenation and QR decomposition and compute the combined weight update:
+Merging these via concatenation and QR decomposition yields the combined weight update.
+Specifically, we first concatenate the basis matrices $B_{l}^{\mathrm{content}}\mid B_{l}^{\mathrm{style}}]$ and apply QR decomposition to their corresponding orthonormal subspaces:
 $Q_{l}^{\mathrm{merged}}=[Q_{l}^{\mathrm{content}}\mid Q_{l}^{\mathrm{style}}]=Q_{l}^{\mathrm{combined}}R_{l}^{\mathrm{combined}}$ (3)
@@
 After rank-limited fine-tuning, we denote the resulting backbone as $W_{\text{init}}=\{W^{\text{init}}_{i}\}_{i=1}^{L}$, which serves as the host weights for all subsequent LoRA training and inference stages.
-3.3.2 Contrastive Content-Style Pairs.
-To facilitate effective disentanglement of content and style during fine-tuning, we construct a carefully curated dataset consisting of $\mathbf{100}$ contrastive content-style image pairs.
-These pairs are specifically designed so that either the content or the style remains constant while the other varies, explicitly guiding the model to differentiate and separately capture content and style characteristics.
+Figure 3: Content-style separation via frequency domain decomposition.
+Each group shows an original image with its frequency-separated elements and additional samples from the same group.
+Left: content group, where low-frequency components capture structural and semantic information.
+Right: style group, where high-frequency components capture textures and artistic rendering.
+Contrastive Content-Style Pairs.
+To facilitate effective disentanglement during fine-tuning, we construct a curated dataset of $\mathbf{100}$ contrastive content-style image pairs.
+These pairs are designed so that either the content or the style remains constant while the other varies, guiding the model to separately capture content and style characteristics.
 We operationalize the notion of content and style using frequency-domain decomposition: content corresponds primarily to low-frequency components encoding structural and semantic information such as object layouts, while style corresponds to high-frequency components capturing textures and color palettes.
 Specifically, we apply Gaussian low-pass filtering with cutoff frequency $\sigma=0.35$ (normalized) to extract content features, and compute the residual as style features.
-These frequency-separated representations are then used as conditioning signals during text-to-image generation to produce the contrastive pairs.
+The cutoff was selected empirically by evaluating separation quality across $\sigma\in\{0.2,0.25,0.3,0.35,0.4,0.45,0.5\}$ via MLLM-based assessment, with $\sigma=0.35$ yielding the best content-style separation across diverse reference images.
 The content and style samples under this frequency method are shown in Figure 3.
-Concretely, we select $\mathbf{10}$ distinct content reference images $\{C_{i}\}_{i=1}^{10}$ and $\mathbf{10}$ distinct style reference images $\{S_{j}\}_{j=1}^{10}$. Exhaustively pairing them yields:
+Concretely, we select $\mathbf{10}$ distinct content reference images $\{C_{i}\}_{i=1}^{10}$ and $\mathbf{10}$ distinct style reference images $\{S_{j}\}_{j=1}^{10}$.
+Exhaustively pairing them yields:
 $\{(C_{i},S_{j})\mid i=1,\dots,10;\;j=1,\dots,10\}.$ (5)
@@
 In this way, the model allocates content and style variations to disjoint low-rank subspaces, ensuring that the backbone initialization $W_{l}^{\mathrm{combined}}$ encodes reduced interference between the two factors.
+We find that diversity matters more than quantity since 50 pairs already achieve 95% of full performance, while increasing to 200 pairs yields less than 2% additional improvement.
 Empirically, combining rank-limited updates with contrastive pairs leads to a 12% reduction in average CLIP-based content–style similarity, indicating that the learned features exhibit reduced cross-influence.
-As further demonstrated in Table 6(b), the removal of rank-limited backbone fine-tuning results in clear drops in both content and style fidelity, quantitatively confirming its role in disentanglement.
+As further demonstrated in Table 6(b), removing Rank-FT results in the largest performance drop among all components, quantitatively confirming its central role in disentanglement.
 Details on dataset construction and training settings are provided in Appendix Appendix.
-Figure 3: Separating content and style components in a single image through frequency domain decomposition
 Algorithm 1 Prompt-Guided Decoupling & Aggregation
@@
 During training, content and style LoRA updates are learned on disjoint layer subsets with explicit prompt markers.
-At inference, an expert encoder routes prompt information to selectively activate these adapters, enabling flexible and fine-grained fusion.
-3.4.1 Prompt-Guided Decoupling and Encoding.
+At inference, an expert encoder routes prompt information to selectively activate these adapters, enabling flexible fusion and fine-grained control.
+Prompt-Guided Decoupling and Encoding.
 Given a raw prompt $p$ containing explicit markers for content and style, e.g.
-p=p=”A photo of a person <c>smiling in a watercolor style <s>”, we treat these markers as routing signals to regulate LoRA module activation and to prevent leakage of content semantics into the style branch and vice versa.
+$p=$ “A photo of a person <c> smiling in a watercolor style <s>”, we treat these markers as routing signals to regulate LoRA module activation and to prevent leakage of content semantics into the style branch and vice versa.
 The prompt with markers removed is encoded by the SDXL text encoder to produce a unified semantic embedding:
@@
 Similarly, we extract content-specific embedding $e_{c}$ and style-specific embedding $e_{s}$ associated with <c> and <s> respectively, enabling isolated semantic guidance for each branch.
-Following the standard LoRA parametrization, each layer $i$ is updated by low-rank matrices.
-To bias the network towards disentanglement, we allocate disjoint index sets $I_{c}$ and $I_{s}$ corresponding to content- and style-sensitive layers: lower and middle layers primarily capture structure and identity (content), while higher layers encode textures and rendering patterns (style).
+Following standard LoRA parametrization, each layer $i$ is updated by low-rank matrices.
+To bias the network towards disentanglement, we allocate disjoint index sets $I_{c}$ and $I_{s}$: lower and middle layers primarily capture structure and identity (content), while higher layers encode textures and rendering patterns (style).
 Content-specific matrices $A_{i}^{(c)}$ and style-specific matrices $A_{i}^{(s)}$ are trained with the semantic embedding, and updates are restricted as
@@
 where $W_{\text{init}}$ denotes the rank-limited backbone obtained in Eqs.3,4.
-The operator $\mathsf{E}_{i}(\cdot)$ injects its input matrix into the slot of the $i$-th layer while filling zeros elsewhere, so the summation is a layerwise direct sum rather than a cross-layer addition.
+The operator $\mathsf{E}_{i}(\cdot)$ injects its input into the $i$-th layer while filling zeros elsewhere, so the summation operates as a layerwise direct sum.
 If markers <c> or <s> are absent, the corresponding branch remains inactive.
-Users may also directly adjust $\gamma_{c},\gamma_{s}$ to control the relative influence of content and style.
-This enables flexible compositions such as “preserve content but replace style,” without retraining.
+Users may also directly adjust $\gamma_{c},\gamma_{s}$ to control the relative influence of content and style, enabling flexible compositions such as “preserve content but replace style” without retraining.
 The full inference procedure is summarized in Algorithm 1, showing how prompts route into disjoint LoRA branches and produce the aggregated backbone $W_{\mathrm{agg}}$.
+Asymmetric Classifier-Free Guidance In standard classifier-free guidance (CFG), the conditional and unconditional paths share the same model weights.
+When LoRA updates are applied, this causes the unconditional path to be contaminated by content and style adapters, leading to instability in image generation.
+To address this issue, we propose Asymmetric Classifier-Free Guidance (ACFG), where the conditional path uses LoRA-adapted weights while the unconditional path remains anchored to the rank-limited backbone $W_{\text{init}}$.
+Let $i\in\{1,\ldots,L\}$ index model layers and $t\in\{1,\ldots,T\}$ diffusion timesteps.
+Instead of uniformly applying LoRA updates at all timesteps, we introduce time-dependent activation schedules $\gamma_{c}(t)=\mathbb{1}_{t\in T_{c}}$ and $\gamma_{s}(t)=\mathbb{1}_{t\in T_{s}}$, where $T_{c}$ and $T_{s}$ denote the active timestep ranges for content and style respectively.
+In practice, we activate content LoRA during early-to-mid timesteps to establish structural layout, and style LoRA during mid-to-late timesteps to refine textures and rendering, reflecting the coarse-to-fine nature of the diffusion process.
+The conditional weights at timestep $t$ are:
+$W^{\mathrm{cond}}(t)=W_{\text{init}}+\sum_{i\in I_{c}}\mathsf{E}_{i}\!\bigl(\gamma_{c}(t)\,\Delta W_{i}^{(c)}\bigr)+\sum_{i\in I_{s}}\mathsf{E}_{i}\!\bigl(\gamma_{s}(t)\,\Delta W_{i}^{(s)}\bigr),$ (9)
+while the unconditional path uses only the initialization backbone throughout:
+$W_{i}^{\mathrm{uncond}}(t)=W^{\text{init}}_{i},\quad\forall\,t,i.$ (10)
+The conditional and unconditional noise predictions are $\epsilon_{\mathrm{cond}}=\epsilon_{\theta}(x_{t}\mid W^{\mathrm{cond}}(t),\mathrm{cond})$ and $\epsilon_{\mathrm{uncond}}=\epsilon_{\theta}(x_{t}\mid W^{\mathrm{uncond}},\varnothing)$ respectively.
+The final guided estimate is:
+$\epsilon_{\theta}^{\mathrm{acfg}}=(1+\omega)\,\epsilon_{\mathrm{cond}}-\omega\,\epsilon_{\mathrm{uncond}},$ (11)
+where $\omega$ controls guidance strength.
+The key insight is that by anchoring the unconditional path to $W_{\text{init}}$, the guidance signal $\epsilon_{\mathrm{cond}}-\epsilon_{\mathrm{uncond}}$ isolates the effect of LoRA adapters at each timestep, rather than conflating adapter influence with the unconditional baseline.
+Combined with the time-dependent schedules, this allows content structure to be established before style details are applied, reducing interference between the two factors.
+Crucially, ACFG requires no additional training or parameter optimization and maintains the same two-pass computation as standard CFG.
+Therefore, ACFG can be directly applied to standard SDXL LoRAs as a single module without the rank-limited backbone, providing improved fusion stability albeit with reduced disentanglement.
 Figure 4: Visual comparison of content-style combinations.
@@
 Competing methods often fail to simultaneously preserve structure and render style, whereas our method produces consistent and coherent content–style compositions.
-3.4.2 Asymmetric Classifier-Free Guidance
-In standard classifier-free guidance (CFG), the conditional and unconditional paths share the same model weights.
-When LoRA updates are applied, this causes the unconditional path to be contaminated by content and style adapters, leading to instability in image generation.
-To address this issue, we propose Asymmetric Classifier-Free Guidance (ACFG), where the conditional path uses LoRA-adapted weights while the unconditional path remains anchored to the rank-limited backbone $W_{\text{init}}$.
-This asymmetry yields a cleaner and more stable guidance signal without requiring additional training.
-Let $i\in\{1,\ldots,L\}$ index model layers and $t\in\{1,\ldots,T\}$ diffusion timesteps.
-Instead of uniformly applying LoRA updates at all timesteps, we introduce time-dependent activation schedules for content and style: $\gamma_{c}(t)=\mathbb{1}_{t\in T_{c}},\gamma_{s}(t)=\mathbb{1}_{t\in T_{s}}.$ The conditional weights at timestep $t$ are defined as
-$W^{\mathrm{cond}}(t)=W_{\text{init}}+\sum_{i\in I_{c}}\mathsf{E}_{i}\!\bigl(\gamma_{c}(t)\,\Delta W_{i}^{(c)}\bigr)+\sum_{i\in I_{s}}\mathsf{E}_{i}\!\bigl(\gamma_{s}(t)\,\Delta W_{i}^{(s)}\bigr),$ (9)
-where $W^{\text{init}}_{i}$ denotes the rank-limited backbone, and $\Delta W_{i}^{(c)},\Delta W_{i}^{(s)}$ are content and style LoRA updates.
-While the unconditional path uses only the initialization backbone:
-$W_{i}^{\mathrm{uncond}}(t)=W^{\text{init}}_{i},\quad\forall\,t,i.$ (10)
-The conditional prediction is $\epsilon_{\mathrm{cond}}=\epsilon_{\theta}(x_{t}\mid W^{\mathrm{cond}}(t),\mathrm{cond})$, and the unconditional prediction is $\epsilon_{\mathrm{uncond}}=\epsilon_{\theta}(x_{t}\mid W^{\mathrm{uncond}},\varnothing)$.
-Then the final guided estimate is
-$\epsilon_{\theta}^{\mathrm{acfg}}=(1+\omega)\,\epsilon_{\mathrm{cond}}-\omega\,\epsilon_{\mathrm{uncond}},$ (11)
-where $\omega$ controls guidance strength.
-By keeping the unconditional path free of LoRA updates, ACFG avoids interference between content and style adapters and provides a more reliable guidance signal.
 Figure 5: Extended visual results.
@@
 “catching a frisbee”, “wearing a hat”, and “driving a car”.
-The results show that our method flexibly integrates dynamic semantics while preserving the specified content and style.
+Our method flexibly integrates dynamic semantics while preserving the specified content and style.
 (b) Single-branch generations where only the content LoRA or only the style LoRA is activated.
-Content-only generations preserve object identity across neutral renderings, while style-only generations transfer artistic features across different subjects, demonstrating disentanglement and controllability.
+Content-only generations preserve object identity across neutral renderings, while style-only generations transfer artistic features across different subjects.
 4 Experiments and Results
 4.1 Implementation Details
-For content concepts, we use the DreamBooth dataset  [27], containing 30 themes with 4-5 images each.
-Style concepts come from the StyleDrop  [33] image set, supplemented with classical and contemporary art style images.
-Importantly, we train each LoRA module (content or style) using only one reference image.
-To enhance content-style separation during backbone fine-tuning, we construct specialized datasets from our main content and style sources.
-These datasets emphasize either content preservation or style transfer properties through a controlled generation process based on frequency separation, as content typically resides in lower frequencies and style in higher frequencies.
-This approach creates clear content-style pairs that provide disentangled signals during backbone fine-tuning, with details provided in the Appendix.Appendix.
-All experiments use the Stable Diffusion XL base model  [25] on NVIDIA 4090 GPUs.
-Our evaluation includes both locally trained LoRAs following our procedure and publicly available pre-trained LoRAs.For locally trained LoRAs, we implement DreamBooth training with rank ($r$) = 64, optimizing weights using Adam-8bit for 1000 steps with batch size 1 and learning rate $1\text{\times}{10}^{-5}$.
-Training uses minimal prompts like ”A [identifier] photo”, where ”[identifier]” (i.e.
-”[c]” for content, ”[s]” for style) serves as a placeholder token representing the specific concept.
+For content concepts, we use the DreamBooth dataset [30] (30 themes, 4-5 images each).
+Style concepts come from StyleDrop [37], supplemented with art style images.
+Each LoRA module is trained from a single reference image.
+We construct 100 contrastive pairs from these sources via frequency-domain decomposition to enhance content-style separation during backbone fine-tuning.
+For evaluation, we use strictly disjoint train/test splits.
+Each metric is averaged over 10 content-style pairs with 10 generated images per pair.
+Content and Style Similarity are computed using CLIP-I embeddings.
+The Combination Score follows MLLM-based binary judgment practices [19]: GPT-4o is presented with the content reference, style reference, and output, and asked whether the output successfully integrates both content identity and artistic style.
+The final score is the average success rate over 50 evaluations per sample.
+[…]
 (a) Quantitative Comparison
@@
 (CLIP-I) $\uparrow$
-Combination Score
+Comb. Score
 (GPT-4o) $\uparrow$
@@
 BLoRA 0.74 0.70 0.77
+LoRA.rar 0.73 0.71 0.76
 Ours 0.79 0.80 0.83
-(b) Ablation study over key components.
-Rank-limited Backbone Fine-tuning
-Setting Content Sim. Style Sim. Cross-Influence Score $\downarrow$
-w/o 0.73 0.69 0.172
-w/ 0.82 0.78 0.087
-Asymmetric Classifier-Free Guidance
-Setting Content Sim. Style Sim. Combination Score $\uparrow$
-w/o 0.77 0.64 0.75
-w/ 0.84 0.78 0.88
+(b) Component ablation study
+Rank-FT Router ACFG Content Sim. $\uparrow$ Style Sim. $\uparrow$ Comb. Score $\uparrow$
+0.65 0.60 0.62
+0.73 0.70 0.72
+0.69 0.65 0.67
+0.68 0.64 0.66
+0.76 0.76 0.80
+0.75 0.73 0.78
+0.71 0.68 0.70
+0.79 0.80 0.83
 (c) User Study
 Method Content Fidelity $\uparrow$ Style Fidelity $\uparrow$ Coherence $\uparrow$
-Direct Merging 45.0% 40.0% 42.5%
-StyleDrop 55.0% 72.5% 60.0%
-ZipLoRA 62.5% 65.0% 67.5%
-KLoRA 65.0% 70.0% 70.0%
-BLoRA 70.0% 67.5% 72.5%
-Ours 82.5% 85.0% 87.5%
+Direct Merging 2.3 2.0 2.1
+StyleDrop 2.8 3.6 3.0
+ZipLoRA 3.1 3.3 3.4
+KLoRA 3.3 3.5 3.5
+BLoRA 3.5 3.4 3.6
+Ours 4.1 4.3 4.4
 Figure 6: Quantitative evaluation of our method.
-(a) Comparison against baselines using CLIP-I similarities and GPT-4o scores.
-(b) Ablation study analyzing the effects of Rank-limited Backbone Fine-tuning and Asymmetric CFG.
-(c) User study results on a 1–5 scale (30 participants, 50 samples), showing clear preference for our method across content fidelity, style fidelity, and overall coherence.
+(a) Comparison against baselines using CLIP-I similarities and GPT-4o binary judgment scores.
+(b) Component ablation: Rank-FT contributes most to disentanglement, Router provides content-style routing, and ACFG stabilizes fusion; all three are complementary.
+The baseline (no components) corresponds to Direct Merging.
+(c) User study ratings on a 1–5 scale (30 participants, 50 samples).
 4.2 Results and Analysis
@@
 Figure 5 provides extended qualitative demonstrations.
-The left part shows content–style combinations augmented with additional prompt descriptions e.g.
+The top part shows content–style combinations augmented with additional prompt descriptions e.g.
 “catching a frisbee”, “playing with a ball”, or “driving a car”.
 Our method integrates these dynamic semantics while faithfully maintaining the designated content and style.
-The right part isolates the effect of single-branch adapters.
+The bottom part isolates the effect of single-branch adapters.
 Activating only the content branch preserves structural fidelity of objects across neutral renderings, while activating only the style branch applies stylistic transformations consistently across varied subjects.
@@
 Each score is averaged over 10 distinct content–style pairs, with 10 generated images per pair.
-As shown in Table 6(a), baseline methods exhibit different trade-offs: Direct Merging fails to preserve either content or style, StyleDrop emphasizes style at the expense of structure, and ZipLoRA, KLoRA, and BLoRA achieve more balanced results but remain limited in overall coherence.
+As shown in Table 6 (a), baseline methods exhibit different trade-offs: Direct Merging fails to preserve either content or style, StyleDrop emphasizes style at the expense of structure, and ZipLoRA, KLoRA, BLoRA, and LoRA.rar achieve more balanced results but remain limited in overall coherence.
 Our method consistently achieves the highest scores across all metrics, demonstrating its ability to preserve structural fidelity while maintaining faithful style transfer and coherent integration.
 Ablation and User Study.
-Table 6(b) analyzes the contributions of our two key components.
-Rank-limited backbone fine-tuning improves CLIP-I content/style similarity and reduces cross-influence, indicating stronger content–style disentanglement.
-The Cross-Influence score measures the relative drop in content/style similarity when composing both factors versus single-factor baselines.
-Specifically, it is computed as the product of CLIP-I similarities between the generated image and both reference images; lower values indicate better disentanglement with reduced mutual interference.
-Asymmetric CFG further strengthens integration quality, reflected by higher human-aligned combination scores and consistent gains in CLIP-I.
-Table 6(c) further reports results from a user study with 30 participants.
-Participants consistently preferred our method across content fidelity, style fidelity, and overall coherence, aligning with the automatic metrics.
+Table 6 (b) presents a comprehensive component ablation.
+Among individual components, Rank-FT contributes most significantly, improving Content Sim.
+by +0.08 and Style Sim.
+by +0.10 over the baseline, confirming its central role in content-style disentanglement.
+The Router provides complementary gains through layer-wise content-style routing, while ACFG stabilizes fusion quality.
+Combining any two components yields further improvements, and the full configuration achieves the best performance across all metrics, demonstrating that the three components are complementary.
+Table 6 (c) reports results from a user study with 30 participants rating 50 samples on a 1–5 scale.
+Our method achieves the highest ratings across content fidelity, style fidelity, and overall coherence, consistent with the automatic metrics.
 This combination of quantitative and perceptual evidence confirms that our framework achieves superior disentanglement, fidelity, and integration compared with prior approaches.
+Failure Cases Analysis.
+Despite the effectiveness of our approach, we identified three main failure modes during the diverse implementation process.
+First, when the reference image contains highly entangled content and style (e.g.
+a cartoon character whose identity is inseparable from the rendering), the method struggles to isolate content from style, resulting in style leakage into the content branch.
+Second, for styles dominated by low-frequency characteristics such as flat color palettes, the frequency-based separation misassigns these features to the content domain, leading to incomplete style capture.
+Third, abstract style descriptions (e.g.
+“ethereal mood”) may lack precise CLIP embeddings, resulting in weak or inconsistent style transfer.
+Adjusting the frequency cutoff $\sigma$ or adopting stronger encoders could mitigate these issues.
 5 Conclusion
 We presented a framework for personalized image generation that integrates efficient adaptation strategies.
-Rank-constrained backbone fine-tuning encourages distinct subspaces, providing a inductive bias for disentanglement.
+Rank-constrained backbone fine-tuning encourages distinct subspaces, providing an inductive bias for disentanglement.
 An expert encoder with selective aggregation enhances semantic guidance, enabling flexible composition, and timestep-dependent asymmetric CFG introduces a training-free sampling strategy that improves visual quality.
-Our experiments substantiate these contributions, demonstrating a measurable reduction in content–style cross-influence, consistent improvements across automatic evaluation metrics, and clear user preference in perceptual studies.
-Limitations include architecture-dependent layer selection, reliance on text embedding quality for semantic extension, and minor overhead from timestep-dependent correction.
-Future work will explore automated layer assignment and extensions to broader diffusion models and modalities.
+Our experiments demonstrate measurable reduction in content-style cross-influence, consistent improvements across automatic metrics, and clear user preference in perceptual studies.
+Limitations include architecture-dependent layer selection, reliance on text embedding quality, frequency-based separation assumptions that may not hold for flat textures, and the two-branch ACFG structure which limits multi-content or multi-style mixing.
+Standard SDXL LoRAs can be used with ACFG directly on the base model, though full disentanglement benefits require retraining on the fine-tuned backbone.
+Future work will explore automated layer assignment, multi-concept scheduling, and extensions to architectures such as KOALA [16] and SANA [42].
 Supplementary Material